How many water molecules are in a bottle of water that contains 500. mL? The density of water is 1.0 g/mL .
1
Expert's answer
2020-12-03T13:53:07-0500

Molar mass of water is 18g.

The bottle has 500 mL of water i.e. 500g of water.

Therefore, no. of moles = 500/18

= 27.78

No. of molecules = 27.78* 6.022*1023

= 1.673 * 1025